# The Riemann Hypothesis is false

### v38

@@
Definitions. Let: $\rho$ denote a non-trivial zero of $\zeta$, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
Lemma 1. (Theorem 4.11 of [7]). We have $\zeta(s)=\sum_{n\leq x}n^{-s}-\frac{x^{1-s}}{1-s}+O(x^{-\sigma})$ uniformly for $\sigma\geq\sigma_{0}>0,|\Im(s)|<2\pi x/C$, when $C$ is a given constant $>1$.
Lemma 2, (Theorem 1 of [3]). One has $\sum_{T\leq\gamma_{T}\leq 2T+1}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ uniformly for $2\leq n\ll T.$
@@
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+\varepsilon+iH}+\int_{1+\frac{1}{\log n}+iH}^{-1+\varepsilon+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{\Theta+\varepsilon+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
@@
say. From now on, let $T\leq t\leq H$ and $\sigma=\Re(s)$. By the approximate functional equation for $\zeta$ (Theorem 4.15 of [7]), note that
@@
Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\Re(s)=\sigma>1$. Inserting this into $I_{1}$ gives
$I_{1}=\frac{(T-H))\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=1,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise). For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{\log(n/m)}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$. Let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=-\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
@@
$\sum_{\frac{T}{\log^{2}T}\leq n\leq\frac{m}{1+\varepsilon}\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (13)
@@
We now treat the ranges $m+1\leq n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq n\leq m-1$. For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=r/m(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=m/r(1+O(r/m))$ hence
@@
Note that if $f(t)$ is differentiable on $[a,b]$, $f^{\prime}(t)$ is monotonic decreasing on $[a,b]$, $f^{\prime}(a)=\beta,f^{\prime}(b)=\alpha,$ then
@@
see e.g. [6, p.10]. Taking $f(u)=\frac{t\log(1+\frac{u}{m})}{2\pi},a=0,b=m\varepsilon-1$ and $f^{\prime}(u)=\frac{t}{2\pi(m+u)}$ gives
@@
$\sum_{1\leq v\leq\alpha}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x=\sum_{1\leq v\leq K_{0}}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $=\sum_{1\leq v\leq K_{0}}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ (25)
@@
Combining (36), (35), (30), (25) and (19), yields $\sum_{a<r\leq b}e^{2\pi if(r)}=o(1)$ and by Abel summation it follows that
@@
For $\frac{m}{1+\varepsilon}\leq n\leq m-1$, put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ thus $\log(n/m)=-r/m(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-m/r(1+O(r/m))$. Thus by arguing as above, one arrives at
@@
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}\ll T^{\Theta}\log T.$ (48)

### v39

@@
Definitions. Let: $\rho$ denote a non-trivial zero of $\zeta,\sigma=\Re(s)$, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
Lemma 1. (Theorem 4.11 of [7]). We have $\zeta(s)=\sum_{n\leq x}n^{-s}-\frac{x^{1-s}}{1-s}+O(x^{-\sigma})$ uniformly for $\sigma\geq\sigma_{0}>0$ and $|\Im(s)|<2\pi x/C$, where $C$ is a given constant $>1$.
Lemma 2, (Theorem 1 of [3]). One has $\sum_{0\leq\gamma\leq y}n^{\rho-1}=-\frac{y\Lambda(n)}{2\pi n}+O((\log y)\log\log 3n)$ uniformly for $2\leq n\ll y.$
@@
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+iH}+\int_{1+\frac{1}{\log n}+iH}^{-1+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{1+\frac{1}{\log n}+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
@@
say. From now on, let $T\leq t\leq H$. By the approximate functional equation for $\zeta$ (Theorem 4.15 of [7]), note that
@@
Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\sigma>1$. Inserting this into $I_{1}$ gives
$I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=1,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise). For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$. Let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=-\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
@@
$\sum_{\frac{T}{\log^{2}T}\leq n\leq\frac{m}{1+\varepsilon}\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (13)
@@
We now treat the ranges $m+1\leq n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq n\leq m-1$. For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=\frac{m}{r}(1+O(r/m))$ hence
@@
Note that if $f(t)$ is a real differentiable function on $[a,b]$, $f^{\prime}(t)$ is monotonic decreasing on $[a,b]$, $f^{\prime}(a)=\beta$ and $f^{\prime}(b)=\alpha,$ then
@@
see e.g. [6, pp.10-11]. Taking $f(r)=\frac{t\log(1+\frac{r}{m})}{2\pi},a=0,b=m\varepsilon-1$ and $f^{\prime}(r)=\frac{t}{2\pi(m+r)}$ gives
@@
$\sum_{1\leq v\leq\alpha}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $=\sum_{1\leq v\leq K_{0}}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ (25)
@@
Combining (34), (33), (29), (24) and (18) yields $\sum_{a<r\leq b}e^{2\pi if(r)}=o(1)$. Thus by Abel summation it follows that
@@
For $\frac{m}{1+\varepsilon}\leq n\leq m-1$, put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ thus $\log(n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-\frac{r}{m}(1+O(r/m))$. Thus by arguing as above, one arrives at
@@
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (46)